Find an expression for the function whose graph consists of the line segment from
the point )2,2 (− to the point )0,1 (− together with the top half of the circle with
centre at the origin and radius 1.
1
Expert's answer
2020-03-02T15:59:50-0500

Find an expression for the function whose graph consists of the line segment from the point (-2, 2) to the point (-1, 0) together with the top half of the circle with center the origin and radius 1

Solution:

To determine:

An expression function for the graph which satisfies the given condition:

Given:

The graph has a line segment connecting (-2,2) and (-1,0) and it consists of a top half of the circle with center (0,0) and radius 1.

Calculation:

Find the slop of the line segment joining the points (-2,2) and (-1,0) as follows:

"m={y_2-y_1 \above{2pt} x_2-x_1}"


"m={0-2 \above{2pt} -1-(-2)}"

"m={-2 \above{2pt} 1}"

"m=-2"

Thus, the slope of the line segment is "m=-2" .

Find the y-intercept of the line segment joining the points (-2,2) and (-1,0) as follows

"y=mx+c"

"0=(-2)(-1)+c" m=-2

"0=2+c"

"c=-2"

Thus , y intercept is c=-2.

The equation of the line segment is "y=-2x-2" for "-2\leqslant x<-1" .

Find the equation of the circle:

The equation of the circle of radius 1 centered on the origin is "x^2+y^2=1"

We can then solve for y

"x^2+y^2=1 \iff" "y^2=1-x^2\iff" "y=\sqrt{1-x^2}"

To only include the top half of the circle we only take the positive root .

The equation of the circle is "\sqrt{x^2-1}" for -1"\leqslant x\leqslant1"

"f(x)=\begin{cases}
 -2x-2 &\text{ on} [-2,1) \\
 \sqrt{1-x^2} &\text{on} [-1,1]
\end{cases}"

Here is a graph;